# Multi-frequency analysis of the ALMA and VLA high resolution continuum observations of the substructured disc around CI Tau

### v1

@@ 2.1 Observations
2.2 (Self-)calibration
ALMA observations –
We followed the standard methodology adopted by the DSHARP collaboration (Andrews et al., 2018), with some minor changes. To begin with, channel-averaging was performed ensuring the same number of channels per SPW in each EB. To avoid bandwidth smearing, we adopted the criterion of Bridle & Schwab (1989) for a reduction of $<1$% in peak response to a point source at the edge of the primary beam. For Band 7, a pseudo-continuum measurement set (MS) was first created, flagging data within $-5\leq v/({\rm km}\,{\rm s}^{-1})\leq 15$ (to account for CI Tau’s systematic velocity) of the 13CO $J=3-2$ ($\nu_{0}=330.588\,{\rm GHz}$), CS $J=7-6$ ($\nu_{0}=342.883\,{\rm GHz}$), and 12CO $J=3-2$ ($\nu_{0}=345.796\,{\rm GHz}$) molecular line transition centres.
When available, the SB data were self-calibrated first, initially in phase-only mode. The gain solutions were computed using the task gaincal [^3] https://casadocs.readthedocs.io/en/v6.4.4/api/tt/casatasks.calibration.gaincal.html. For the first run, this was done separately for each polarisation (gaintype=‘G’) for Band 3 and combining both polarisations (gaintype=‘T’) for Band 6, combining different scans, with an infinite solution interval. From the second run on, for both Band 3 and 6 data, the gain solutions were computed with gaintype=‘T’, combining different scans and spectral windows, with solution intervals progressively decreasing from 240, 120, 60, 30, to $15\,{\rm s}$. Images were reconstructed after each iteration using the tclean task. We set the CLEANing threshold to three times the noise level, used 10 pixels per beam, a multi-scale multi-frequency synthesis deconvolver (mtmfs, see Rau & Cornwell 2011) with nterms=2, and adopted a briggs weighting scheme (Briggs, 1995) with robust=0.5. We estimated an improvement of the peak signal-to-noise ratio (SNR) of $\approx 28$% and $\approx 40$% for Band 3 and Band 6 data, respectively. Finally, a round of phase and amplitude self-calibration was performed. The gain solutions were computed with gaintype=‘T’, combining different scans and spectral windows, with an infinite solution interval. However, we did not observe any improvement of the peak SNR (as already reported by Konishi et al. 2018 for Band 6 data).
Then, all the EBs in the same band were concatenated and self-calibrated together. To remove spatial offsets, we measured the emission centroid positions of each dataset with a Gaussian fit in the image plane. Subsequently, the phaseshift task was used to move the disc centre to the fitted phase centre position, and the fixplanets task to align each EB to the phase centre of the LBs with the highest peak SNR (i.e., the EB with the smallest astrometric uncertainty and lowest noise). Before these steps, we performed a run of phase-only self-calibration on the LB data to increase their peak SNR ratio. The gain solutions were computed with gaintype=‘G’, combining different scans with an infinite solution interval. This additional step helped improving the alignment procedure. Subsequently, the visibilities of each dataset were deprojected and azimuthally averaged. To do so, we used the median of the posterior distribution of the inclination ($i=49.24\,{\rm deg}$) and position angle (${\rm PA}=11.28\,{\rm deg}$) that Clarke et al. (2018) determined by fitting the Band 6 LB data in the Fourier space parametrically with galario (Tazzari et al., 2018). These deprojected visibilities were then inspected to identify and correct for mismatches in the amplitude scales. This step is crucial to ensure an accurate flux scaling with wavelength, and proved to be particularly important for Band 6 data, where flux density variations larger than 10% over a single day were measured. We refer to Appendix A for a detailed discussion of our correction procedure, and the time variability of the flux calibrators.
The short and long baseline data in the same ALMA Bands were then combined using the concat task. We performed a number of phase-only self-calibration iterations on the concatenated MS. For the first run, the gain solutions were computed with gaintype=‘G’ for Band 3 and 7, and gaintype=‘T’ for Band 6, combining different scans with an infinite solution interval. From the second run on, the gain solutions were computed with gaintype=‘T’, combining different scans and spectral windows, with solution intervals progressively decreasing from 360, 180, to 60 for Band 3, and to 30 and $15\,{\rm s}$ for Band 6 and 7. Image reconstruction was performed as for the SBs, but with robust=1.0. We estimated an improvement of the peak SNR of $\approx 21$% for Band 6, $\approx 26$% for Band 7, and only marginal for Band 3. Finally, we performed a round of phase and amplitude self-calibration with gaintype=‘T’, combining different scans and spectral windows with an infinite solution interval. We flagged the most extreme gain solutions (with gains $<0.8$ or $>1.2$), that are associated with long-baseline antennas and artificially reduced the beam size. After this iteration, the peak SNR ratio improved substantially only in Band 7 by an additional factor of $\approx 20$% (i.e., a total improvement by 56%).
VLA observations –
We conducted spectral averaging on each dataset, considering the necessary precautions to prevent bandwidth smearing [^4] https://casadocs.readthedocs.io/en/v6.3.0/notebooks/synthesis_imaging.html. Time averaging was not performed. In the imaging process, we adopted a briggs weighting scheme with robust=1.0, representing the optimal trade-off between SNR and side lobe effects. The mtmfs deconvolver with nterms=2 was employed to accurately account for the large bandwidth-over-observing-frequency ratio in the VLA data.
For the Ka Band observations, we initially aligned the data from different EBs to ensure that the disc centre was consistent across all the datasets. Then, we applied a single round of phase-only self-calibration only to the B configuration (lower angular resolution) data, leading to a slight improvement of the peak SNR by $\approx 5\%$. We adopted the following gaincal parameters: gaintype=’G’, combine=’scan’, solint=’inf’. Subsequently, the B configuration data were combined with the A configuration (higher angular resolution) ones. Similarly, a round of phase-only self-calibration was performed, resulting in a further increase in the peak SNR by $\approx 10\%$. Additional self-calibration iterations with shorter solution intervals did not yield any further improvement. For the Ku Band data, we applied one round of phase-only self-calibration (again with gaintype=’G’, combine=’scan’, solint=’inf’) resulting in a marginal improvement of the SNR by $\approx 10\%$. Due to the initially (too) low SNR, no self-calibration was performed on the X and C Band data to avoid introducing spurious effects.
@@
Band Frequency Robust Beam size Beam PA RMS noise Peak intensity Peak SNR Flux density $R_{68}$ $R_{95}$
@@
Band 7 338.228 $-1.0$ $0.056\times 0.086$ $-33.20$ $1.05\times 10^{-1}$ 12.55 119.48 $410.16\pm 0.82$ $0.77_{-0.01}^{+0.01}$ $1.22_{-0.01}^{+0.01}$
$-0.5$ $0.059\times 0.088$ $-33.09$ $7.30\times 10^{-2}$ 13.02 178.27 $413.80\pm 0.60$ $0.78_{-0.01}^{+0.01}$ $1.24_{-0.01}^{+0.01}$
0.0 $0.065\times 0.096$ $-32.94$ $5.15\times 10^{-2}$ 14.32 278.00 $417.51\pm 0.82$ $0.79_{-0.01}^{+0.01}$ $1.26_{-0.01}^{+0.02}$
0.5 $0.077\times 0.111$ $-33.69$ $4.14\times 10^{-2}$ 16.77 404.76 $416.02\pm 0.69$ $0.79_{-0.01}^{+0.01}$ $1.27_{-0.01}^{+0.01}$
1.0 $0.092\times 0.133$ $-34.67$ $3.95\times 10^{-2}$ 19.94 505.03 $418.27\pm 0.91$ $0.81_{-0.01}^{+0.01}$ $1.28_{-0.01}^{+0.01}$
1.5 $0.100\times 0.145$ $-34.90$ $4.12\times 10^{-2}$ 21.71 526.90 $418.98\pm 1.36$ $0.81_{-0.01}^{+0.01}$ $1.29_{-0.01}^{+0.01}$
Band 6 229.612 $-1.0$ $0.021\times 0.033$ 7.95 $2.89\times 10^{-2}$ 1.73 59.74 $146.88\pm 0.89$ $0.71_{-0.04}^{+0.03}$ $1.17_{-0.02}^{+0.01}$
$-0.5$ $0.022\times 0.034$ 7.94 $2.05\times 10^{-2}$ 1.80 88.09 $146.90\pm 0.57$ $0.71_{-0.02}^{+0.02}$ $1.18_{-0.01}^{+0.01}$
@@
Band 3 97.504 $-1.0$ $0.034\times 0.050$ $-11.22$ $1.44\times 10^{-2}$ 0.60 41.51 $15.85\pm 0.07$ $0.57_{-0.06}^{+0.04}$ $1.09_{-0.02}^{+0.01}$
$-0.5$ $0.037\times 0.054$ $-10.25$ $1.04\times 10^{-2}$ 0.66 63.36 $16.20\pm 0.07$ $0.59_{-0.04}^{+0.04}$ $1.10_{-0.01}^{+0.01}$
0.0 $0.044\times 0.064$ $-6.22$ $7.55\times 10^{-3}$ 0.79 104.85 $16.42\pm 0.08$ $0.62_{-0.03}^{+0.04}$ $1.15_{-0.02}^{+0.01}$
@@
Ka Band 33.0 $-1.0$ $0.042\times 0.045$ $-66.48$ $5.07\times 10^{-3}$ 0.18 34.86 $0.32\pm 0.06$ $<0.04$ $0.05_{-0.01}^{+0.01}$
$-0.5$ $0.046\times 0.053$ $-47.58$ $3.37\times 10^{-3}$ 0.18 52.86 $0.32\pm 0.03$ $<0.05$ $0.07_{-0.01}^{+0.01}$
0.0 $0.057\times 0.068$ $-45.25$ $2.39\times 10^{-3}$ 0.18 76.05 $0.33\pm 0.02$ $<0.06$ $0.08_{-0.01}^{+0.01}$
0.5 $0.086\times 0.100$ $-49.15$ $1.77\times 10^{-3}$ 0.19 105.83 $0.31\pm 0.02$ $<0.09$ $0.17_{-0.01}^{+0.01}$

### v2

@@ Appendix I Absorption spectral index
Band Frequency Robust Beam size Beam PA RMS noise Peak intensity Peak S/N Flux density $R_{68}$ $R_{95}$
@@
Band 7 338.228 1.0-1.0 $0.056\times 0.086$ 33.20-33.20 $1.05\times 10^{-1}$ 12.55 119.48 $410.16\pm 0.82$ $0.77_{-0.01}^{+0.01}$ $1.22_{-0.01}^{+0.01}$
0.5-0.5 $0.059\times 0.088$ 33.09-33.09 $7.30\times 10^{-2}$ 13.02 178.27 $413.80\pm 0.60$ $0.78_{-0.01}^{+0.01}$ $1.24_{-0.01}^{+0.01}$
0.0 $0.065\times 0.096$ 32.94-32.94 $5.15\times 10^{-2}$ 14.32 278.00 $417.51\pm 0.82$ $0.79_{-0.01}^{+0.01}$ $1.26_{-0.01}^{+0.02}$
0.5 $0.077\times 0.111$ 33.69-33.69 $4.14\times 10^{-2}$ 16.77 404.76 $416.02\pm 0.69$ $0.79_{-0.01}^{+0.01}$ $1.27_{-0.01}^{+0.01}$
1.0 $0.092\times 0.133$ 34.67-34.67 $3.95\times 10^{-2}$ 19.94 505.03 $418.27\pm 0.91$ $0.81_{-0.01}^{+0.01}$ $1.28_{-0.01}^{+0.01}$
1.5 $0.100\times 0.145$ 34.90-34.90 $4.12\times 10^{-2}$ 21.71 526.90 $418.98\pm 1.36$ $0.81_{-0.01}^{+0.01}$ $1.29_{-0.01}^{+0.01}$
Band 6 229.612 1.0-1.0 $0.021\times 0.033$ 7.95 $2.89\times 10^{-2}$ 1.73 59.74 $146.88\pm 0.89$ $0.71_{-0.04}^{+0.03}$ $1.17_{-0.02}^{+0.01}$
0.5-0.5 $0.022\times 0.034$ 7.94 $2.05\times 10^{-2}$ 1.80 88.09 $146.90\pm 0.57$ $0.71_{-0.02}^{+0.02}$ $1.18_{-0.01}^{+0.01}$
@@
Band 3 97.504 1.0-1.0 $0.034\times 0.050$ 11.22-11.22 $1.44\times 10^{-2}$ 0.60 41.51 $15.85\pm 0.07$ $0.57_{-0.06}^{+0.04}$ $1.09_{-0.02}^{+0.01}$
0.5-0.5 $0.037\times 0.054$ 10.25-10.25 $1.04\times 10^{-2}$ 0.66 63.36 $16.20\pm 0.07$ $0.59_{-0.04}^{+0.04}$ $1.10_{-0.01}^{+0.01}$
0.0 $0.044\times 0.064$ 6.22-6.22 $7.55\times 10^{-3}$ 0.79 104.85 $16.42\pm 0.08$ $0.62_{-0.03}^{+0.04}$ $1.15_{-0.02}^{+0.01}$
@@
Ka band 33.0 1.0-1.0 $0.042\times 0.045$ 66.48-66.48 $5.07\times 10^{-3}$ 0.18 34.86 $32\pm 0.06 <0.0$4<0.04 $0.05_{-0.01}^{+0.01}$
0.5-0.5 $0.046\times 0.053$ 47.58-47.58 $3.37\times 10^{-3}$ 0.18 52.86 $32\pm 0.03 <0.0$5<0.05 $0.07_{-0.01}^{+0.01}$
0.0 $0.057\times 0.068$ 45.25-45.25 $2.39\times 10^{-3}$ 0.18 76.05 $33\pm 0.02 <0.0$6<0.06 $0.08_{-0.01}^{+0.01}$
0.5 $0.086\times 0.100$ 49.15-49.15 $1.77\times 10^{-3}$ 0.19 105.83 $31\pm 0.02 <0.0$9<0.09 $0.17_{-0.01}^{+0.01}$